OpenStudy (anonymous):

Krissy has 3 arrows. The probability of her hitting the target is 2/5. What is the probability that she will get at least one arrow on the target?

OpenStudy (mathmate):

At least one means she does not miss all, which is easier to calculate. Probability of missing all three is P(0)=(3/5)^3 so P(1+)=1-P(0)

OpenStudy (anonymous):

What would the answer be in percent?

OpenStudy (mathmate):

P(0)=(3/5)^3= 0.216 so P(1+)=1-P(0)=0.784

OpenStudy (anonymous):

Thank you!

OpenStudy (mathmate):

You're welcome!